# A microscopic theory of Anderson localization of electrons in random lattices

### v1

@@ IV.2.2 Localized phase
The localized phase is more difficult and less straightforward to describe. The reason is that the micrscopic irreducible vertex $\Lambda_{0}(E;\omega)=\Lambda_{R}(E;\omega)+i\Lambda_{I}(E;\omega)$ has a discontinuity in its imaginary part at zero frequency, $\Lambda_{I}(E;0_{\pm})=\pm|\Lambda_{I}(E;0)|$. This jump is transferred to the conserving vertex and the electron-hole correlation function. Its denominator in the localized phase in the mean-field approximation has a more complex structure than assumed in Ref. [1] because of a nontrivial structure of the bound electron-hole pair
@@
$\mathcal{D}(E;\omega,\mathbf{q})\doteq\frac{1+i\alpha_{0}(E)}{2|\Im\Sigma^{R}(E)|}\\ \times\left[-i\left(\omega+\frac{2\alpha_{0}(E)|\Im\Sigma^{R}(E)|}{1+\alpha_{0}(E)^{2}}\right)+\ D(E)q^{2}\right.\\ \left.-\ \frac{2\alpha_{0}(E)^{2}|\Im\Sigma^{R}(E)|}{1+\alpha_{0}(E)^{2}}\right]$ (68)
@@
where we denoted $\alpha_{0}(E)=\Lambda_{I}(E;0)/\Lambda_{R}(E;0)$ and used the mean-field diffusion constant
@@
The terms proportional to $\alpha_{0}$ cause the breakdown of the linear response theory, and Eq. (58) for the dynamical conductivity cannot be used. Instead, the extension of the time propagation resulting from the electron-hole correlation function should be analyzed at long distances, in the limit $q\to 0$. It is easy to see that there is no long-range diffusion if
@@
Diffusion can be restored if the particle gets sufficient energy to break the bond of the electron-hole pair. The threshold wave vector $q_{0}$ above which diffusion is restored in the localized phase then is
@@
where $m$ is the mass of the particle. The threshold energy can be acquired by the applied electric field $U_{b}=E_{b}/|e|$. The electric current will flow for electric fields $U>U_{b}$, and the system becomes metallic.
@@ V Conclusions
There are various options and effective statistical models for simulating and studying the critical behavior of the Anderson localization transition. It is, however, necessary to analyze the model of propagation of a quantum particle in a random environment to understand the microscopic origin of the vanishing of charge diffusion and the emergence of Anderson localization. The fundamental quantities to be studied in the thermodynamic limit are configurationally averaged one-particle and two-particle resolvents. The microscopic quantities fully characterizing them are the one-particle self-energy and the two-particle irreducible vertices. The Anderson localization transition and the phases on both sides of the transition must be evidenced by these functions.
@@
We presented a general scheme for constructing approximations in renormalized perturbation theory that interpolate between the weak- and strong-disorder limits, including the Anderson localization transition. There are three principal aspects to make the approximations successful. First, the two-particle vertex must contain more information about the impact of the disorder on the particle propagation than provided by the one-particle self-energy. Second, the microscopic dynamical quantities must be separated from the macroscopic observables that obey conservation laws, since causality of the self-energy and the two-particle vertices is incompatible with the dynamical Ward identities. Third, local and nonlocal scattering must be treated separately, as the two-particle self-consistency required for Anderson localization can only be achieved with nonlocal, off-diagonal one-particle propagators.
@@
Perturbation theory uses two-particle-irreducible vertices as its fundamental ingredients to provide a more detailed picture than the self-energy. The consistency between the one-particle and two-particle Green functions is achieved by matching the imaginary part of the self-energy with the real part of the static and homogeneous electron-hole irreducible vertex via the Ward identity. The complete dynamical Ward identity cannot be demanded in the approximations for the dynamical two-particle vertices. The conserving two-particle vertex is restored from the dynamical one at the end via a given relation so that the conservation laws of the response functions and macroscopic observables are obeyed.
@@
The perturbation theory for the two-particle irreducible vertices must be split into the local and nonlocal contributions. The local contributions are summed via the coherent potential approximation. Its local vertex serves as the input for the expansion beyond this solution to obtain nontrivial two-particle vertex functions. Nonlocal scattering allows one to distinguish between electron-electron (cooperon) and electron-hole (diffuson) multiple scattering. Their self-consistent interconnection yields two-channel parquet equations that are nonlinear integral equations with a bifurcation critical point for strong disorder, representing the Anderson localization transition.
@@
We employed a local approximation to the integral kernel of the parquet equations to obtain quantitative results for the behavior of the conserving electron-hole vertex, the electron-hole correlation function, and the dynamical conductivity in both metallic and localized phases. We received a mean-field critical behavior at the Anderson localization transition and an order parameter in the localized phase. We calculated mean-field characteristics of an electron propagating in the localized phase, spatially restricted by a quantum bound state with a hole left at the origin. We assessed the localization length and the energy barrier that must be overcome to drive an electric current.
@@
Our construction provides a microscopic explanation of Anderson localization, the emergence of a quantum bound state between a propagating particle and the hole left at the origin. The existence of a bound state is signalled by a gap in the dynamical conductivity, corresponding to a threshold energy required to break the bound state, impose an electric current, and turn the electron’s localized wave function into an extended one. This result can be generalized to conclude that a gap in a two-particle response function, without a gap in the self-energy, indicates the existence of quantum bound states without their condensation. Linear response theory breaks down beyond the localization transition, and static quantities with no memory, such as conductivity and the diffusion constant, cannot describe particle propagation in the Anderson-localized phase. Only the fully dynamical response functions contain imprints of Anderson localization.

### v4

@@ IV.2.3 Localized phase
The behavior in the localized phase is more difficult to describe and less straightforward. The perturbation expansion in small frequency breaks down. The reason for this is that the microscopic irreducible vertex $\Lambda_{0}(E;\omega)=\Lambda_{R}(E;\omega)+i\Lambda_{I}(E;\omega)$ has a discontinuity in its imaginary part at zero frequency, $\Lambda_{I}(E;0_{\pm})=\pm|\Lambda_{I}(E;0)|$. This jump is transferred to the conserving vertex and the electron-hole correlation function. In the localized phase, its denominator in the mean-field approximation has a more complex structure than assumed in Ref. [1] because of the nontrivial structure of the bound electron-hole pair
@@
$\mathcal{D}(E;\omega,\mathbf{q})\doteq\frac{1+i\alpha_{0}(E)}{2|\Im\Sigma^{R}(E)|}\\ \times\left[-i\left(\omega+\frac{2\alpha_{0}(E)|\Im\Sigma^{R}(E)|}{1+\alpha_{0}(E)^{2}}\right)+\ D(E)q^{2}\right.\\ \left.-\ \frac{2\alpha_{0}(E)^{2}|\Im\Sigma^{R}(E)|}{1+\alpha_{0}(E)^{2}}\right]\,$, (71)
@@
with $\alpha_{0}(E)=\Lambda_{I}(E;0)/\Lambda_{R}(E;0)$. We used the mean-field diffusion constant
@@
The terms proportional to $\alpha_{0}$ cause the breakdown of linear response theory, and Eq. (58) for the dynamical conductivity cannot be used. Instead, the time propagation in the electron-hole correlation function should be analyzed at long distances, in the limit $q\to 0$. It is easy to see that there is no long-range diffusion if
@@
Diffusion can be restored if the particle acquires sufficient energy to break the bond of the electron-hole pair. The threshold wave vector $q_{0}$ above which diffusion is restored in the localized phase is then
@@
where $m$ is the particle’s mass. The threshold energy is given by the applied electric field $U_{b}=E_{b}/|e|$. The electric current flows in the direction of the electric field when $U>U_{b}$, and the system becomes metallic.
@@ V Conclusions
There are various options and effective statistical models for simulating and studying the critical behavior of the Anderson localization transition. However, to understand the microscopic origin of the vanishing of charge diffusion and the emergence of Anderson localization, it is necessary to analyze the original model of a quantum particle propagating in a random environment. The fundamental quantities to be studied in the thermodynamic limit are the configurationally averaged one-particle and two-particle resolvents. The microscopic quantities fully characterizing them are the one-particle self-energy and the two-particle irreducible vertices. These functions must evidence the Anderson localization transition and the phases on both sides of the transition.
@@
We presented a general scheme for constructing approximations in renormalized perturbation theory that interpolate between the weak- and strong-disorder limits, including the Anderson localization transition. There are three principal aspects that make the approximations successful. First, the two-particle vertex must capture more of the disorder’s impact on particle propagation than the one-particle self-energy. Second, the microscopic dynamical quantities must be separated from the macroscopic observables that obey conservation laws, because causality of the self-energy and the two-particle vertices is incompatible with the dynamical Ward identities. Third, local and nonlocal scattering processes must be treated separately, as the two-particle self-consistency required for Anderson localization can be achieved only with nonlocal, off-diagonal one-particle propagators.
@@
Perturbation theory uses two-particle-irreducible vertices as its fundamental ingredients to provide a more detailed picture than the self-energy. The consistency between the one-particle and two-particle Green functions is ensured by matching the imaginary part of the self-energy with the real part of the static and homogeneous electron-hole irreducible vertex via a Ward identity, Eq. (25). The complete dynamical Ward identity cannot be imposed on approximations for the dynamical two-particle vertices. The irreducible conserving two-particle vertex, $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$, is recovered from the dynamical one, $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$, via a given relation, Eqs. (46), (47), ensuring that the conservation laws of the response functions and macroscopic observables are satisfied.
@@
The contributions to the two-particle irreducible vertices must be split into local and nonlocal terms. The local contributions are summed using the coherent potential approximation. Its local vertex serves as input for the expansion beyond this solution to obtain nontrivial two-particle vertex functions. Nonlocal scattering allows one to distinguish between electron-electron (cooperon) and electron-hole (diffuson) multiple scattering. Their self-consistent interconnection yields two-channel parquet equations, nonlinear integral equations with a bifurcation critical point for strong disorder, representing the Anderson localization transition.
@@
We employed a local, mean-field approximation to the parquet equations’ integral kernel to obtain quantitative results for the behavior of the conserving electron-hole vertex, the electron-hole correlation function, and the dynamical conductivity in both metallic and localized phases. We observed a mean-field critical behavior at the Anderson localization transition and the genesis of an order parameter in the localized phase. We calculated mean-field characteristics of an electron propagating in the localized phase, spatially restricted by a quantum bound state with a hole left at the origin. We assessed the localization length and the energy barrier required to drive an electric current in the localized phase.
@@
The presented construction provides a microscopic explanation of Anderson localization as the emergence of a quantum bound state between the propagating particle and the hole left behind. The existence of a bound state is signaled by a gap in the dynamical conductivity, corresponding to the threshold energy required to generate an electric current and turn the electron’s localized wave function into an extended one. This result can be generalized to conclude that a gap in a two-particle response function, without a gap in the self-energy, indicates the existence of quantum bound states without their condensation. Linear response theory breaks down beyond the localization transition, and static quantities with no memory, such as conductivity and the diffusion constant, cannot describe particle propagation in the Anderson-localized phase. Only the fully dynamical response functions contain imprints of Anderson localization.
We conclude that, in this paper, we have completed the construction of a consistent mean-field theory of a new phase in disordered electron systems, extending Anderson localization to dynamical quantities. We found that the new phase emerges bekow an instability line in the $(E,\omega)$-plane, where the frequency derivative of the homogeneous electron-hole correlation function, $\partial\Im\Phi^{RA}(E;\omega,\mathbf{0})/\partial\omega$ from Eq. (8), changes sign from positive above the line (normal metallic phase) to negative below it (semimetallic phase). The best quantity to check this behavior is the electric polarizability.